# Moving towards informative and actionable social media research

### v1

@@
Joseph B. Bak-Coleman 1,2,3,4,5∗ Stephan Lewandowsky 6,7 Philipp Lorenz-Spreen 8,9∗ Arvind Narayanan 10,11 Amy Orben 12 Lisa Oswald8
@@ 12 MRC Cognition and Brain Sciences Unit, University of Cambridge, Cambridge, UK
Social media is nearly ubiquitous in modern life, and concerns have been raised about its putative societal impacts, ranging from undermining mental health and exacerbating polarization to fomenting violence and disrupting democracy. Despite extensive research, consensus on these effects remains elusive, with observational studies often highlighting concerns while randomized controlled trials (RCTs) yield conflicting or null findings. This review examines how the complexity inherent in social systems can account for such discrepancies, emphasizing that emergent societal and long-term outcomes cannot be readily inferred from individual-level effects. In complex systems, such as social networks, feedback loops, hysteresis, multi-scale dynamics, and non-linearity limit the utility of approaches for assessing causality that are otherwise robust in simpler contexts. Revisiting large-scale experiments, we explore how null or conflicting findings may reflect these complexities rather than a true absence of effects. Even in cases where the methods are appropriate, assessing the net impacts of social media provides little actionable insight given that eliminating social media is not a realistic option for whole populations. We argue that progress will require a complexity-minded approach focused on specific design choices of online platforms that triangulates experimental, observational and theoretical methods.
@@ Introduction
Social media have permeated most societies, impacting citizens lives around the world. Nearly 5 billion people across the globe are using social media, and penetration rates exceeding 80% in northern Europe, 70% in North-, 67% in South-America and reaching 75% in Eastern Asia, for example  (?). The defining attribute of social media platforms is that users not only consume but also create and distribute content. In addition, social media are characterized by some degree of algorithmic curation and in most cases also an advertisement-driven business model  (?).
@@
Accompanying the pervasive use of social media is widespread concern about their adverse effects on society, ranging from undermining public health measures  (?) and mental health  (?, ?) to exacerbating polarization  (?), fomenting extremism  (?), and disrupting democracy  (?). A plethora of research has been seeking to address those concerns; Google Scholar returns more than 800,000 publications since 2020 in response to the query “social media” (Search on 16 February 2024). At least tacitly, this substantial research endeavor is concerned with generating a consensus causal understanding of the relationships between social media use and the outcomes of interest. Absent such an understanding, any necessary corrective action (including national or international regulation and intervention) may remain ineffective.
Unfortunately, most efforts trying to understand the impact of social media on large-scale societal changes —to ultimately judge whether they warrant corrective action— persist in the liminal space between compelling patterns or individual case studies of harm on one hand, and consensus causal estimates on the other. For example, a recent systematic review of the literature of social media’s impact on political behavior identified almost 500 papers  (?) and revealed a multi-dimensionality of associations. Many reviewed papers reported positive associations between social-media use and political participation and knowledge, and at least as many papers reported associations between social media use and growing polarization and success of populist parties, which point in a detrimental direction for liberal democracies.
Since then, most researcher agree to the correlational observations that were reported by Lorenz-Spreen et al., no strong consensus on the causal impacts and mechanisms of social media on society and democracy has emerged. Scientists remain divided about whether polarization is caused, amplified or just made visible by social media platforms  (?) and how or even whether that causes harm  (?). Similar discussions are also had regarding the impact of social media on mental health, violence and other societal outcomes of interest. Specifically, due to the lack of agreement on causal mechanisms, there are few (if any) concrete suggestions for policies or regulations that should be implemented to reduce harms and promote positive individual and societal outcomes in our digital age.
The inability to reach scientific consensus and provide actionable insight hints at the existence of deeper problems in the research endeavor that require resolution before progress can occur. Here, we argue that there are two key barriers to an effective science of social media. First, social media constitute complex systems that comprise ensembles of individual elements that interact with each other in specific ways, thereby forming higher-level order and emergent phenomena on the collective level (?, ?). However, scientists tend to use techniques such as randomized controlled trials (RCTs) that are well suited to simpler, socially-independent individual outcomes but ill-suited to establishing causality for complex systems outcomes (e.g., societal polarization), particularly in the context of social media. This leads to long-winded debate, as bias, confounds, and model assumption failures may drive observed effects more than the underlying phenomena of interest—a general challenge recently reviewed in the social sciences more broadly by Bailey et al  (?). Second, even when empirical circumstances allow answers to certain causal questions, various studies tend to produce only broad insights (e.g., about average effects of deactivation) that are not directly actionable by technologists and regulators.
To overcome those barriers, we propose (1) a complex systems approach to study social media’s impacts on society, which is based on the triangulation of mechanistic models, experiments suitable for estimating causal effects at a relevant scale, as well as descriptive work and observational causal inferences at levels of systemic complexity. (2) Furthermore, we suggest to focus research questions on more detailed understanding of relationships between platform affordances and behaviors or attitudes.
Research on Social Media’s Impacts
@@
Research on social media’s impacts routinely uncovers associations suggesting links between platform use and broader societal or individual outcomes. For example, consistent correlational evidence has been identified across several dimensions, including between social media use and political participation, polarization, trust in political institutions, body image and well-being  (?, ?, ?). Aiming to gain causal insight into these associations, some studies have used quasi-experimental instrumental-variable approaches  (?) in which predictable variation of local internet adoption or social media use caused by a third variable (the “instrument”) has been used to estimate the causal impact on trust  (?, ?), political participation  (?, ?), hate crimes  (?, ?), or populist support and polarization  (?, ?), as well as mental health  (?, ?). The results from those naturalistic quasi-experiments yield a relatively consistent and concerning pattern, with social media being shown to increase hate crimes, lower trust, or increase polarization and support for populists (for a more detailed overview, see  (?)).
An alternative approach involves large-scale field experiments in which individuals are placed into randomized controlled trials (RCTs). Treatments are often constrained by limitations in independent researchers’ abilities to alter platform experiences and functionality. Examples range from participants following or subscribing to political accounts on Twitter/X and Facebook  (?, ?) through deactivation of Facebook for a month or a week  (?, ?, ?) to setting specific news websites as the starting page of a browser  (?), browsing allocated websites of a specific valence  (?) or implementing a suite of different approaches to reduce digital distractions  (?). More recent studies, conducted in collaboration with Meta, were able to go beyond these typical limitations, actively altering news-feed algorithms  (?, ?, ?). Another, recent study used a browser add-on to reduce toxic content within the feed and without direct collaboration with the platforms  (?).
Notably, results from those RCTs rarely seem to produce consensus in the social media context, (?) with the resulting findings conflicting both across RCTs as well as with inferences drawn from observational or quasi-experimental approaches. For example, two studies in which the treatment involved subscriptions/exposure to specific content affected polarization in opposite directions  (?, ?). A similar divergent pattern was observed between the two deactivation experiments of Facebook  (?, ?), where one study found reduced polarization from the deactivation the other found increased animosity against outgroups from the same treatment. In some cases, RCTs yield effects that are consistent with concerns raised from observational causality—yet effect sizes are small and may fail to achieve statistical significance. For example, in the most recent and largest deactivation experiment of Facebook and Instagram, affective polarization was slightly reduced and voting choice appeared altered  (?). However, this finding was judged not to be significant after application of a correction for multiple comparisons.
To further complicate the matter, outcomes from RCTs can differ substantially depending on what is measured. Results from interventions are often the most conflicting or weak when examining distal outcomes, such as polarization. In contrast, the impact of specific interventions on more proximal outcomes, such as exposure to harmful content, appear to be more stable. For example, the Meta collaboration generated consistent evidence that the type of content that individual people are exposed to is impacted by platform design and affordances, such as platform-mediated amplification of co-partisan or uncivil content  (?, ?, ?). On the other hand, the same studies showed that those changes had virtually no downstream consequences on outcomes of particular interest such as political attitudes  (?, ?, ?). As a result, there seems to be a missing link, whereby RCTs find consistent evidence with observational approaches for proximal outcome measures on the level of individual content exposure—yet effects on people’s distal behaviors and attitudes are conflicting or minimal. This creates a challenge for those who seek to attribute causal responsibility to platform algorithms for distal or societal-level outcomes such as polarization.
We argue that this apparent gap between individual-level outcomes (e.g., content exposure, engagement, sharing behaviour) and societal outcomes (e.g., polarization, out-group attitudes) arises because social media are inherently complex systems in which production and consumption of content are entangled in feedback loops that often operate across different time scales and levels of organizational complexity. Failing to account for such complexity when designing RCTs may compromise a causal interpretation of the results. Further, by casting aside these crucial attributes of complex systems, researchers have frequently focused on research questions that are in principle unanswerable or unactionable. If these questions are instead approached within a complex-systems perspective and associated methods, we will be one step closer to research that can lead to actionable policy, regulation, or design changes.
Causality in complex systems
There are certain questions about social media that are no longer answerable by direct observation. Just as we will never know what world we might inhabit now had the automobile not become the primary means of transport in America, we will never know how politics might have unfolded in the absence of Facebook or Twitter. The inability to examine counterfactual realities is known as the fundamental problem of causal inference, and is a widespread challenge across scientific research  (?).
Although causality cannot be directly measured in a strict sense, this has not stopped scientists from developing techniques to infer it indirectly  (?). Historically, the RCT is the principal workhorse in establishing causality. In an RCT, individual units of interest are randomly assigned to groups (e.g., treatment and control), and differences between those groups are evaluated. This design sidesteps the fundamental problem of causal inference by assuming that randomization, at least in expectation, eliminates systematic differences between the groups and yields an unbiased impact of treatment (but see  (?)). Early in their training, scientists are often taught that RCTs are the “gold-standard” of causality, which frequently results in misunderstanding of precisely how RCTs work, what they require, and what they yield  (?).
Most conspicuously, individual-level RCTs cannot reliably yield the estimand of interest for most complex social phenomenon, such as the causal effect of social media on collective processes such as polarization, extremism, mental health trends or vaccine uptake. To do so, one would have to randomize social media exposure at the scale of societies or (ideally) worlds. In lieu of this, RCTs are typically applied at the scale of individual users, under the buried assumption that societal effects over longer periods of time (which is often the actual estimand of interest) are equal to additive individual effects over much shorter periods of time (the estimated quantity). This assumption that causal impacts at societal scales should be observable at the individual level is tenuous, specifically for studying the societal impact of social media, which is a quintessential example of a complex system.
Complex systems are defined by the way in which interactions lead to emergent phenomena such that the behavior of the system cannot be understood as the sum of its parts (cf., ”More is Different”  (?)). Moreover, complex systems exhibit dynamics such that any estimated quantity is not independent of the point in time and historical state. Non-linearities introduced by such interactions produce feedback, phase-transitions, and hysteresis which preclude the notion of some time-invariant causal effect.
@@
We highlight four key ways in which RCTs may yield misleading interpretations of causal effects both for effects that occur on longer time-scales or at larger scales of organizational complexity (see Figure 1.) We can illustrate those problems using an intuitive physical example. Consider concerns from observational evidence that power-grid blackouts coincide with increased demand for power in a world where the relationship between power demand and system failure is not known. Attempting to test this theory, an RCT is conducted on a power grid in which some substations are randomized to experience sudden spikes in demand. If the treatment group receiving increased load is small relative to the scale of the grid, the power grid will remain functional as the grid compensates for the localized demand. One would be forced to conclude that spikes of that magnitude have no effect on the power grid’s functionality. Conversely, if a critical number of substations had been exposed to the same spike in demand, blackouts would occur—including the control nodes for which no increase in demand occurred, because the blackouts would affect both types of nodes equally. Again, this would result in little difference between the groups, suggesting that blackouts occur independent of the condition (spike or not). Both of those interpretations would be incorrect because the estimated quantity from an individually targeted RCT is not the estimand of interest (also see Fig. 1A). In each of these cases, the outcome of interest cannot be isolated to individuals as the impacts of the treatment alter the system as a whole, in turn altering outcomes for both treatment and control groups alike.
The power grid example can also provide insight into the second problem, hysteresis (Figure 1B), which occurs when the relationship between inputs and the system state is not a simple function but depends on the past state of the system. Following failure to a power grid, an RCT relieving the load for some substations, but not others, would fail to bring any stations back online. Withdrawal studies rest on the assumption that the effect of the platform or feature can be inferred as the opposite of the measured withdrawal effect. We would be wrong to conclude from such a study that load did not cause the failure, merely because eliminating the load did not return function. This problem of interpretation besets all withdrawal studies, in which users in the treatment group are deactivating a platform, reducing use or even removing a feature  (?, ?, ?, ?).
In social systems, “path-dependence” is commonly used to describe hysteretic effects. To illustrate, suppose we want to evaluate the causal effect of Duolingo on language fluency, and we do so by randomly selecting long-time users to discontinue learning for a week. We are bound to find that users largely retain their fluency after a week of withdrawal of the platform. Of course, it would be incorrect to assume the lack of change in fluency reflects an inability of Duolingo to cause language learning. In general, there is no reason to believe that changes in beliefs, behaviors, attitudes (e.g., trust in institutions), or alterations of mental health acquired over years of time spent on a social media platform will rapidly or symmetrically revert upon brief (e.g., weeks) withdrawal, if such reversals are even possible or likely (also see Fig. 1B). Similar considerations apply to other social outcomes: For example, mental health might transition from one stable state (healthy) to another (unhealthy), but going back is not simply an additive feature of “undoing” the original trigger  (?). In general, we cannot reasonably assume that withdrawal from a social media platform produces equal and opposite effects as the initial exposure.
@@
The third problem associated with interpreting individual-level RCTs is that the effects may unfold non-monotonically over time and depend on both the state of other variables in the system and the moment of observation (Figure 1C). For example, the feedback loop of addiction means that having a cigarette can lead to a relief (e.g., reduced tension or stress) in a long-term smoker. When such long-term smokers are asked to withdraw from smoking for two weeks, they are likely to become irritable (and gain weight, among other side effects of withdrawal). The clear health benefits of smoking cessation would only be observable over much larger time-scales. Under the reasoning common in social media studies RCTs, short-term studies of smoking cessation would incorrectly conclude that smoking is a safe and effective means of losing weight and reducing irritability (also see Fig. 1C). Likewise, relevant outcomes on social media may exhibit temporal dynamism—specifically recommendation algorithms that learn from user behaviour (?)—so we should anticipate that findings across experiments of differing lengths will conflict and short-term results may be of very limited value to extrapolate to the long term.
@@
Finally, perhaps the core reason why RCTs are an inappropriate tool to study macro-outcomes on social media and power grids alike is that in both cases the individual units interact (Figure 1D). Formally, the problem here is that any RCT would violate the “stable unit treatment value assumption” (SUTVA), which requires that the treatment outcome of one experimental unit (e.g., a participant or power substation) is not affected by the treatment status of other units  (?). The violation of SUTVA impacts our powergrid RCT above, because whether or not a substation can continue running depends not only on its local load, but on the load which other directly and indirectly connected stations are experiencing and the power they are producing.
@@
This problem also affects RCTs on social media. People are coupled through a myriad of on- and off-platform connections. An individual randomly exposed to an experimentally-altered newsfeed, for example by switching to a chronological sorting, will nonetheless experience algorithmic curation: When neighbors in their network share content they saw through algorithmic curation, the individual in the chronological condition is still indirectly subject to algorithmic sorting. Similarly, the mental well-being of a teenager taken off social media will not just depend on that experimental manipulation, but on whether their friends or parents also stop using these platforms.
@@
Ultimately, violations of SUTVA can be avoided only if entire subnetworks are randomized into treatment and control. That is, all friends and family on social media must be assigned to the same treatment (or control) as the focal person of interest whose behaviour is recorded. To our knowledge, no existing RCTs fulfill that criterion.
In summary, any individually-focused RCT on social media that is not explicitly addressing the problems just presented—i.e., SUTVA violations, hysteresis, feedback loops and non-linearity inherent to complexity—is bound to be of limited value as clear causal identification of macro-level outcomes is not possible. The estimated magnitude and even direction of effects may be misleading, and there is no simple way to correct or adjust for such biases post hoc. It follows that any assertion that RCTs are interpretable or superior to observational approaches simply by virtue of being RCTs is not sustainable. The quality of any causal insight depends wholly on the assumptions underlying the causal inference and not the intrinsic properties of the approach. Very often the identification of a causal mechanism requires a multitude of methods, for example by triangulating it from observational, experimental and theoretical angles  (?). We explore examples of this problem using several recent studies in depth in the next section.
Example: RCTs on Facebook
@@
A recent collaboration between Meta and independent researchers attempted to bridge the gap in our causal understanding of political outcomes algorithmic curation, using RCTs. In the experiments, some users were exposed to a number of interventions, such as receiving less politically-aligned content in their news feeds  (?), experiencing a reduced number of direct reshares from social contacts  (?), being switched to a chronologically sorted feed  (?), or deactivating the entire platform for 6 weeks  (?). Generally speaking, these interventions exhibited effects on exposure to content. For example, default platform features (e.g., algorithmic sorting) tended to expose users to more content that was ideologically aligned and uncivil but also less moderate sources, political content and untrustworthy sources compared to a reverse-chronological feed  (?).
@@
At first glance, these findings are consistent with some concerns that engagement-optimized social media exacerbates polarization and foments hate via increased exposure to uncivil and like-minded content. However, the studies additionally conducted surveys to evaluate downstream effects on relevant attitudes, beliefs and behaviors. Here, effects were generally smaller, less consistent, and more ambiguous. For example, reducing exposure to content from like-minded sources did not decrease affective polarization or ideological extremity  (?). The authors asserted their estimated causal effects on individuals provided insight into the societal-scale estimand  (?).
@@
We can reappraise those studies in light of the issues surrounding complexity raised above. First, as with our power grid and increased energy demand, there is little reason to believe that changes in collective state (e.g., polarization) from increased platform use or platform-feature exposure can be directly inferred from changes in exposure for an ensemble of individuals. Such changes involve not only the direct effects of exposure (or withdrawal), but also the changes following exposure to the dynamics across individuals (global outcomes, see Fig. 1A). Moreover, much like a power station remaining dark once demand is reduced, platform-polarized individuals may not revert to a more moderate state merely because exposure to the platform or content is reduced (hysteresis, see Fig. 1B). This applies not just to individuals, but likewise to the structure of the network. In the case of polarization, for example, the network of co-partisan individuals one follows could be heavily shaped by algorithmic recommendations, yet these changes are not undone by temporarily down-ranking co-partisan content in one’s own feed (Spillover, see Fig. 1D).

### v2

@@
Joseph B. Bak-Coleman 1,2,3,4,5,6∗ Stephan Lewandowsky 7,8 Philipp Lorenz-Spreen 9,10∗ Arvind Narayanan 11,12 Amy Orben 13 Lisa Oswald10
@@ 13 MRC Cognition and Brain Sciences Unit, University of Cambridge, Cambridge, UK
Social media is nearly ubiquitous in modern life, raising concerns about its societal impacts—from mental health and polarization to violence and democratic disruption. Yet research on its causal effects remains inconclusive: observational studies often find concerning associations, while randomized controlled trials (RCTs) tend to yield small, conflicting, or null results. Literature summaries tend to causally prioritize findings from RCTs, often arguing that concerns about social media are overstated. However, like observational studies, RCTs rely on assumptions that can easily be violated in the context of social media, especially regarding societal outcomes at scale. Here, we enumerate and examine the features of social media as a complex system that challenge our ability to infer causality at societal scales. Drawing on insight from disciplines that have faced similar challenges, like climate-science or epidemiology, we propose a path forward that combines the strength of observational and experimental approaches while acknowledging the limitations of each.
@@ Introduction
Social media have permeated most societies, impacting citizens’ lives around the world. Nearly 5 billion people across the globe are using social media, and penetration rates exceed 80% in northern Europe, 70% in North-, 67% in South-America, and reach 75% in Eastern Asia (?). The defining attribute of social media platforms is that users not only consume but also create and distribute content. In addition, social media are characterized by some degree of algorithmic curation and in most cases also an advertisement-driven business model (?).
@@
Accompanying the prominence of social media is widespread concern about their adverse effects on society, ranging from undermining public health measures (?) and mental health (?, ?) to exacerbating polarization (?), fomenting extremism (?), and disrupting democracy (?). A plethora of research has sought to address those concerns; Google Scholar returns more than 800,000 publications since 2020 in response to the query “social media” (Search on 16 February 2024). At least tacitly, this substantial research endeavor is concerned with generating a consensus causal understanding of the relationships between social media use and the outcomes of interest. Absent such an understanding, any necessary corrective action, from design to regulation, may remain ineffective.
@@
Unfortunately, most efforts trying to understand the impact of social media on large-scale societal changes — to ultimately judge whether they warrant corrective action — are compelling enough to raise concern but fail to yield a scientific consensus. Broadly speaking, observational evidence has consistently uncovered associations between social media use and political participation, polarization, trust in political institutions, body image, and well-being (?, ?, ?). Aiming to gain causal insight into these associations, some observational work has leveraged quasi-experimental instrumental-variable approaches (?) to estimate the causal impact on trust (?, ?), political participation (?, ?), hate crimes (?, ?), or populist support and polarization (?, ?), as well as mental health (?, ?). The results from those naturalistic quasi-experiments yield a relatively consistent and concerning pattern, with social media being shown to increase hate crimes, lower trust, or increase polarization and support for populists (for a more detailed overview, see (?)).
Yet experimental findings often reach the opposite conclusion—finding small, null and heterogeneous effect sizes. Often, these involve Randomized Controlled Trials (RCTs) that modify some aspect of users’ experience with social media. A wide range of such experiments have been carried out, ranging from following specific accounts or installing experience-altering browser extensions to news-feed modification or abstention from platforms altogether (?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?). For societally-relevant outcomes such as polarization, extremism, and voting patterns, the estimates that arise from this body of literature tend to be the least clear (?).
Often the resolution to conflicting observational and experimental work has been to assume or assert that experiments provide better, less-biased causal estimates. For example, high-profile experiments conducted in conjunction with Meta were argued to “rule out” moderate sized or larger effects of their platforms on various societal outcomes such as extremism and polarization (?, ?, ?). Similarly, a review of the impacts of misinformation cited experiments as providing causal evidence which ostensibly trumped concerns stemming from observations, qualitative work, and journalism (?). In the context of adolescent mental health, an authoritative report by the National Academies declared that longitudinal evidence could not be used to establish causality, implying formal experiments were sufficient and necessary (?).
In this article, we critically examine those claims. We show that they arise from twin misunderstandings that (a) RCTs are sufficient and necessary for establishing causality and that (b) observational research cannot support causal claims (?, ?). We highlight how a defining feature of social media—social interactions, often facilitated through algorithms—can produce all manner of feedback, emergent system-level properties, and temporal dynamics (see Fig. 1 for an overview) (?, ?) that undermine the assumptions required by RCTs and likely leads to downward-biased effect sizes—a general challenge recently reviewed in the social sciences more broadly by Bailey et al. (?). Additionally, we discuss how, when, and whether the estimates generated by RCTs meaningfully correspond to the causal effects on society.
However, even if we resolve long-standing debates about social media’s net historical impact on society, it may leave us far from the kind of scientific insight needed to guide intervention by policymakers and technologists. To overcome these barriers, we draw on insights from other scientific disciplines, such as climate change and public health, where feedbacks and interactions also frustrate the standalone use of tools such as RCTs. Following in their footsteps, we highlight a path forward for social media research that embraces the strengths of a wide variety of approaches while appreciating and offsetting their limitations. Beyond benefits for uncovering mechanism, we argue this approach will enable research to focus on questions that reveal more mechanistic, actionable insight into the relationships between platform design, use, and societal phenomena.
Causally Inferring Social Media’s Effects
We cannot simultaneously observe the presence and absence of a treatment on a single unit (e.g., person, society, etc.). There is no way to directly measure whether an individual or society would be better or worse off with social media. This is true in science more broadly and known as the “fundamental problem of causal inference” (?). To work around this problem, scientists instead rely on relating some source of variation to an outcome of interest. For observational studies, the source of variation is naturally occurring, whereas experiments artificially induce variation in the exposure to treatment. Often, this variation is induced through randomization into treatment and control (i.e., an RCT), artificially varying exposure.
Observational approaches to inferring causality have proliferated in recent decades, following the development of techniques such as Directed Acyclic Graphs (DAGs) and do-calculus (?, ?). These approaches involve formally defining putative causal relationships between variables of interest and, in theory, enable causal identification by testing assumptions about the relationships between ostensible causes and effects. For instance, natural variation in the roll-out of Facebook across college campuses was used as source of variation to draw inferences about its impacts on mental health (?). Similarly, variation in smartphone policies in schools have been used to examine impacts on educational and mental health outcomes (?). Outages and feature roll-outs provide additional sources of variation to be exploited. Often, these sources of variation are framed as “instrumental variables” which impact treatment exposure but are not believed to impact outcomes.
To illustrate, natural disasters divert media attention from any other topics they would normally focus on, including violent events such as mass shootings. This can be exploited by comparing the downstream consequences of mass shootings when media coverage was or was not coincidentally suppressed by a natural disaster overseas. Such comparison reveals that domestic media coverage of U.S. mass shootings causes further mass shootings: a one standard deviation increase in shooting news raises mass shootings by approximately 73% of a standard deviation, supporting the notion of behavioral contagion when mass shootings are rendered salient by media coverage (?). Because no plausible causal pathway exists between natural disasters taking place abroad and mass shootings in the U.S., the association between media coverage and further mass shootings can be considered causal.
Any of these observational approaches to causality require specific assumptions to hold for unbiased causal estimation. Non-random roll-outs or application of smartphone bans could, for instance, confound their treatment effect. Addressing these challenges requires adding co-variates to the statistical model to adjust for sources of bias and ideally identify the causal pathway of interest (?, ?, ?).
To avoid these complexities, scientist routinely opt to instead induce variation by assigning treatments to units, often at random. At least in expectation, the randomization into treatments and controls addresses confounding and yields an unbiased causal estimate (?). Yet this feature often leads to a misunderstanding that RCTs intrinsically yield causal inference, when they too draw causal inference from assumptions about the nature of, and relationships between, measured values (?). For example, differential attrition and chance bias introduced during randomization are two of the more commonly considered pitfalls, leading authors to estimate conditional treatment effects (e.g., Intent to Treat, Local Average Treatment Effect) or include a variety of co-variates to the statistical model.
Societal vs. individual causal effects
Yet social media research poses challenges for identifying causal effects using RCTs that are largely unappreciated in the literature. First and foremost, it is essential that any estimate of a causal effect is indeed the causal effect of interest (i.e., the estimand). Many areas of concern regard phenomena that occur at societal scales, such as polarization, extremism, trends in mental health, or democratic outcomes. By societal scales, we mean that the phenomenon as they occur across groups of individuals. For example, while mental health can be studied as an individual phenomenon, its trends and changes over time are implicitly societal phenomena.
@@
Distinctions of scale are rare in prominent experiments on social media’s impacts, which often take the form of randomizing individuals to induce variation (?, ?, ?). Yet in doing so, the estimand implicitly changes from the effect on societal phenomena to the effect on individuals. Taking the estimate from such a study to indicate a societal effect requires an often unstated assumption—that societal effects over longer periods of time (which is often the estimand of interest) are equal to the simple sum of individual effect on a shorter time scales (the estimated quantity, Fig. 1A). Claims about the past effect of a platform on some outcome additionally require assumptions that the direction and magnitude of an effect is time and scale-invariant; from the birth of the platform with its first user to the time-frame of the experiment.
Unfortunately, causal effects on society will rarely be identical to the effects on individuals. Social media is defined by interactions between individuals, a hallmark of complex systems more broadly. Although the term “complex” takes on many meanings across contexts, here we are referring to systems in which interactions between individual parts yield emergent system-level properties (?, ?). A defining feature of emergent properties in complex systems is that they cannot be inferred from the simple sum of the system’s constituent parts (cf.,“More is Different” (?, ?)) (Fig. 1A).
As a concrete example, a nation-wide experiment on Facebook involved a treatment whereby users could share that they had voted and see how many others had done the same (?). The direct, individual treatment effect yielded 60,000 additional votes compared to a control group which did not see social information. However, the authors went one step further and examined social contagion effects, which led to an additional 280,000 votes. In this case, the societal estimand is more than five times larger than the individual estimand. Positive-feedback, such as the amplification of voter turnout, is a common feature of complex systems and can frustrate estimation of societal effects from individual outcomes. Negative-feedback, by contrast, can dampen such effects or even potentially reverse their direction. Often, the societal effect will depend on a mix of positive and negative feedback, making it very difficult to infer the societal effect from individual effects alone.
Hysteresis
Even estimating the causal effect on individuals is a challenging task on social media platforms. The ubiquity of social media makes it challenging, if not impossible, to find a representative sample of non-users. Owing to this and potential ethical concerns, withdrawal experiments reducing access to social media platforms or platform features are often used (?, ?, ?, ?). However, treatment effects from withdrawal experiments are implicitly confounded. Past use of the social media platform or feature is required for treatment assignment, such that past use exerts indirect causal effects through the treatment assignment to the outcome of interest. Randomization is intended to capture this causal effect of treatment. Yet if such causal effect exists, it also exerted an effect prior to the study on the outcome of interest, creating a classic confound.
To illustrate, suppose we want to evaluate the causal effect of Duolingo (a language learning platform) on language fluency, and we do so by randomly selecting long-time existing users to discontinue learning for a week. We are bound to find that users largely retain their fluency after a week of withdrawal of the platform. Of course, it would be incorrect to assume the minimal decline in fluency reflects minimal causal effect of Duolingo on language skills accumulated over years of learning. Here, the confound of past use has resulted in downward bias of the estimated effect size. Moreover, for anything that is learned faster (or slower) than it is acquired the estimate will be further biased accordingly.
In general, there is no reason to believe that changes in beliefs, behaviors, attitudes (e.g., trust in institutions), or alterations of mental health acquired over years of time spent on a social media platform will rapidly or symmetrically revert upon brief (e.g., weeks or months) withdrawal, if such reversals are even possible or likely (also see Fig. 1B).
Similar considerations apply to other social outcomes: For example, mental health might transition from one stable state (healthy) to another (unhealthy) because of social media, but reverting to a healthy state is not simply an additive feature of “undoing” the original trigger (?). This lack of reliable symmetry between the effect of a cause and its reversion is a common feature of complex systems known as hysteresis. Just as we might not expect individuals to return to a prior state, there is little reason to expect that societal effects of social media will revert as platform use, features, and affordances change. It is conceivable, for instance, that a change in societal phenomena becomes self-reinforcing such that the platform’s causal effect is akin to a match on dry tinder. Once the fire is burning, the effect of the match would become difficult to identify. As such, hysteresis presents challenges for both observational and experimental research alike.
Feedback Loops
@@
The third problem associated with interpreting individual-level RCTs is that the effects of interest may unfold non-monotonically over time (Figure 1C). For example, the feedback loop of addiction means that having a cigarette can lead to a relief (e.g., reduced tension or stress) in a long-term smoker. When such long-term smokers are asked to withdraw from smoking for two weeks, they are likely to become irritable (and gain weight, among other side effects of withdrawal). The clear health benefits of smoking cessation would only be observable over much larger time-scales. Under the reasoning common in social media studies involving RCTs, short-term studies of smoking cessation would incorrectly conclude that smoking is a safe and effective means of losing weight and reducing irritability (also see Fig. 1C). Likewise, relevant outcomes on social media may exhibit temporal dynamism—specifically recommendation algorithms that learn from user behaviour (?)—so we should anticipate that findings across experiments of differing lengths will conflict and short-term results may be of very limited value to extrapolate to the long term.
Spillover/SUTVA
@@
Finally, perhaps the core reason why RCTs are an inappropriate tool to study macro-outcomes on social media is that in both cases the individual units interact (Figure 1D). Formally, the problem here is that any RCT would violate a dimension of the “stable unit treatment value assumption” (SUTVA)—or non interference—which requires that the treatment outcome of one experimental unit (e.g., a participant) is not affected by the treatment status of other units (?).
@@
People are coupled through a myriad of on- and off-platform connections. An individual randomly exposed to an experimentally-altered newsfeed, for example by switching to a chronological sorting, will nonetheless experience algorithmic curation: When neighbors in their network share content they saw through algorithmic curation, the individual in the chronological condition is still indirectly subject to algorithmic sorting. Similarly, the mental well-being of a teenager taken off social media will not just depend on that experimental manipulation, but on whether their friends or parents also stop using these platforms. It follows that violations of SUTVA can be generally expected to downward bias any results, deflating the true magnitude.
@@
Ultimately, violations of SUTVA can be avoided only if entire subnetworks are randomized into treatment and control. That is, all friends and family on social media must be assigned to the same treatment (or control) as the focal person of interest whose behaviour is recorded. To our knowledge, no existing RCTs studying societal effects of social media fulfill that criterion. Far from hypothetical, downward bias from SUTVA has been a known issue to social media companies for decades and strategies have been developed to address it (e.g., network randomization) in their own A/B testing of product features (?). Curiously, industry-academic collaborations studying societal effects nonetheless relied on individual randomization (?, ?).
A common theme emerges from the challenges of applying RCTs to infer causal societal effects. Small numbers of individual effects can nonetheless lead to large societal phenomena, as demonstrated in the example of voting behavior on Facebook described above. Hysteresis too, would be anticipated to generally produce downward bias—especially when past use is a confound. Positive feedback loops would similarly be anticipated to yield experimental under-estimation when effect are measured on shorter time-scales than the feedback takes to manifest. Finally, spillover and SUTVA violations will tend towards systematic underestimation on social media (?).
Although overestimation is certainly possible, the general expectation from the features above would be systematic downward bias (i.e., towards zero) of effect sizes. We provide an example of overestimation later in the piece, whereby the effect of misinformation interventions is likely over-estimated. Nevertheless, the frequent null and near-zero effects observed in social media studies may often be an artifact of downward bias rather than true absence of meaningful effects. Unfortunately, as many of these issues are implicit to study design, there is no universal way to post hoc adjust for such bias.
Example: RCTs on Facebook and Instagram
@@
A recent collaboration between Meta and independent researchers attempted to bridge the gap in our causal understanding of political outcomes algorithmic curation, using RCTs. In the experiments, some users were exposed to a number of interventions, such as receiving less politically-aligned content in their news feeds (?), experiencing a reduced number of direct reshares from social contacts (?), being switched to a chronologically sorted feed (?), or deactivating the entire platform for 6 weeks (?). Generally speaking, these interventions altered users’ exposure to content. For example, default platform features (e.g., algorithmic sorting) tended to expose users to more content that was ideologically aligned and uncivil but also less moderate sources, political content and untrustworthy sources compared to a reverse-chronological feed (?).
@@
At first glance, these findings are consistent with some concerns that engagement-optimized social media exacerbates polarization and foments hate via increased exposure to uncivil and like-minded content. However, the studies additionally conducted surveys to evaluate downstream effects on relevant attitudes, beliefs and behaviors. Here, effects were generally smaller, less consistent, and more ambiguous. For example, reducing exposure to content from like-minded sources did not decrease affective polarization or ideological extremity (?). The authors asserted that their estimated (null) causal effects on individuals provided insight into the societal-scale estimand (?).
@@
We can reappraise those studies in light of the issues surrounding complexity raised above. First, there is little reason to believe that changes in collective state (e.g., polarization) from increased platform use or platform-feature exposure can be directly inferred from changes in exposure for an ensemble of individuals. Such changes involve not only the direct effects of exposure (or withdrawal), but also the changes following exposure to the dynamics across individuals (global outcomes, see Fig. 1A). Moreover, platform-polarized individuals may not revert to a more moderate state merely because exposure to the platform or content is reduced (hysteresis, see Fig. 1B). This applies not just to individuals, but likewise to the structure of the network. In the case of polarization, for example, the network of co-partisan individuals one follows could be heavily shaped by algorithmic recommendations, yet these changes are not undone by temporarily down-ranking co-partisan content in one’s own feed (Spillover, see Fig. 1D).